OpenStudy (anonymous):

Find the length of a side of a square whose diagonal is 4 cm

OpenStudy (amistre64):

what do we know about the properties of a square that would help us use the pythag thrm here?

OpenStudy (anonymous):

c^2 = a^2 * b^2 c=diagonal, a and b are the sides. But since this is a square, a and b are equal.

OpenStudy (anonymous):

Now do a little algebra...

OpenStudy (amistre64):

a^2 + b^2 that is :)

OpenStudy (anonymous):